### v1

# On the Federated Learning Framework for Cooperative Perception

Abstract
Cooperative perception is essential to enhance the efficiency and safety of future transportation systems, requiring extensive data sharing among vehicles on the road, which raises significant privacy concerns. Federated learning offers a promising solution by enabling data privacy-preserving collaborative enhancements in perception, decision-making, and planning among connected and autonomous vehicles (CAVs). However, federated learning is impeded by significant challenges arising from data heterogeneity across diverse clients, potentially diminishing model accuracy and prolonging convergence periods. This study introduces a specialized federated learning framework for CP, termed the federated dynamic weighted aggregation (FedDWA) algorithm, facilitated by dynamic adjusting loss (DALoss) function. This framework employs dynamic client weighting to direct model convergence and integrates a novel loss function that utilizes Kullback-Leibler divergence (KLD) to counteract the detrimental effects of non-independently and identically distributed (Non-IID) and unbalanced data. Utilizing the BEV transformer as the primary model, our rigorous testing on the OpenV2V dataset, augmented with FedBEVT data, demonstrates significant improvements in the average intersection over union (IoU). These results highlight the substantial potential of our federated learning framework to address data heterogeneity challenges in CP, thereby enhancing the accuracy of environmental perception models and facilitating more robust and efficient collaborative learning solutions in the transportation sector.
Index Terms:
Cooperative intelligent transportation system, cooperative perception, autonomous driving, federated learning, bird’s-eye-view segmentation.
I Introduction
Environmental perception is crucial for advanced driver assistance systems (ADAS) and autonomous driving, encompassing tasks like instance segmentation [2, 3, 1, 4], traffic scene understanding [5], 2D object detection [6], 3D object detection [7, 8, 9, 10], and multi-object tracking [11, 12, 13, 14, 15]. These technologies also play a critical role in enhancing cooperative intelligent transportation systems (C-ITS), significantly improving transportation safety and efficiency [16, 17, 18]. The rise of C-ITS has drawn considerable interest from both academia and industry, highlighting a collaborative framework that integrates various ITS components—personal, vehicle, roadside, and central systems—to surpass traditional ITS models in service quality. Connected automated vehicles (CAVs) exchange information through vehicle-to-everything (V2X) networks, including vehicle-to-vehicle (V2V), vehicle-to-infrastructure (V2I), vehicle-to-network (V2N), and infrastructure-to-network (I2N).
A major challenge within C-ITS is the security and privacy of shared data. Federated learning in general offers a robust solution by enabling decentralized model training without compromising data privacy, which has already been demonstrated in many applications like wireless communications [19], Industry 4.0. [20] and robotics [21], etc. In C-ITS, this approach also allows CAVs to collaboratively participate in environment perception, traffic flow prediction, and decision-making, to improve deep neural network (DNN) models while keeping the privacy protected by avoiding exchange sensor data or latent features between CAVs [22]. Additionally, V2X communication supports seamless data exchange between vehicles, roadside units (RSUs), and cloud systems, enhancing the effectiveness of federated learning in C-ITS [23, 24, 25]. Supported by RSUs, digital twins can produce accurate labels for learning algorithms, thereby increasing the precision and performance of the systems.
Despite its numerous advantages, the federated learning paradigm encounters a significant challenge in data heterogeneity when implemented across diverse clients. This issue is particularly pronounced in C-ITS perception which is also known as cooperative perception (CP) due to the varied sensor configurations among participating vehicles and infrastructure. Such variations include discrepancies in camera counts, sensor placements, and non-independently and identically distributed (Non-IID) data among traffic agents, leading to marked differences in data characteristics [26]. This diversity presents a significant challenge for traditional federated learning methods, as it may be challenging to effectively synchronize the contributions of local models towards the convergence of the global model. This paper explores strategies to harness the benefits of external data while mitigating the impacts of data heterogeneity within a federated learning framework, with the aim of significantly improving the performance of local models.
Among the critical tasks essential for CP, bird’s-eye-view (BEV) perception is emphasized due to its key role in achieving comprehensive environmental awareness. To address the challenge of data heterogeneity, we introduce a novel federated learning framework specifically designed for BEV perception. Specifically, we proposed an advanced aggregation methodology within our federated learning framework, to confront with varied data distributions among clients and their disparate impacts on global model convergence. This aggregation mechanism is designed to handle the heterogeneity intrinsic to diverse agents, thereby enhancing training efficiency and model efficacy. Additionally, we introduced a novel loss function designed to mitigate unstable convergence issues, steering the model toward alignment with the global optimizer.
To validate our approach with fair comparison, we conducted extensive experiments and follow the same BEV vision model proposed in [27], which consists of five main components: an encoder for image feature extraction; positional embedding to capture camera geometry; a cross-attention module that transitions front views to BEV; convolution layers within transformers for feature refinement; and a decoder to convert BEV representations into actionable predictions. The results demonstrate the superiority of our framework, showing improved test accuracy and reduced communication overhead across most clients, thus confirming the effectiveness of the proposed federated learning framework in addressing the complex challenges.
Our contributions are summarized as follows:
  • A federated learning framework tailored for CP task with a novel federated dynamic weighted aggregation (FedDWA) algorithm is introduced. This framework addresses data heterogeneity across clients by employing a dynamic constant mechanism, which adjusts the balance between global and local model, narrowing the performance gap and fostering a cohesive learning environment.
  • To enhance training efficacy further, a dynamic adjusting loss (DALoss) function is proposed. This function is tailored to the varying data distributions among clients and the central server, fine-tuning the model’s convergence direction based on real-time data distribution insights, thus ensuring improved convergence and significantly boosting the accuracy of BEV perception outcomes.
The rest of the paper are organised as follows: Section II provides related work, discussing CP, federated learning, and the application of federated learning within the CP context. Section III details our federated learning algorithms: It begins with a formulation of the problem and describes the optimization objectives of our federated learning framework, then our proposed FedDWA algorithm and DALoss function is introduced. In Section IV, our experimental setup and datasets are explained. Our experimental results, including the comparison between our proposed algorithm with other state-of-the-art methods and corresponding ablation study are analyzed. Finally, the conclusion is drawn in Section V.
II Related Work
II-A Cooperative Perception in Autonomous Driving
It is essential that autonomous driving systems have the ability to perceive the environment accurately. This perception directly affects the efficiency of decision-making and overall safety. Traditional systems, constrained by limited sensing capabilities, often struggle with occlusions and distant objects. CP, enabled by V2X communications, marks a significant improvement, allowing CAVs to enhance perception through extensive wireless information sharing, encompassing V2V, V2I, and V2N communications. OpenCDA, introduced by [28], provides a versatile framework for developing and testing cooperative driving automation (CDA) systems within a simulated environment. Complementarily, the work [29] presented a large-scale simulated dataset for V2V perception, enriched with a diverse collection of scenes and annotated data. This dataset proves critical for advancing CP systems, as supported by [16], which underscores the expanded sensing capabilities facilitated by V2X. Further developments include the CoBEVT framework introduced in [30], a sparse vision transformer tailored for autonomous driving that leverages multi-agent multi-camera sensors, enabling collaborative perception across multiple vehicles. The work [31] discusses the integral role of advanced algorithms in optimizing CP systems. The works [32] and [33] explore the potential of integrating infrastructure-based sensors to bolster CP. There have been some notable advancements in perception frameworks recently. One of these is the introduction of V2X-ViT [34]. V2X-ViT is a novel vision Transformer designed to effectively combine information from on-road agents. Another important contribution is the work by Xu et al. [35]. They have created the first extensive real-world multimodal dataset for V2V perception. This dataset significantly enhances data availability for CP research. Lastly, Xiang et al. have introduced HM-ViT in their paper [36]. HM-ViT is an innovative multi-agent hetero-modal cooperative perception framework. It can accurately predict 3D objects in dynamic V2V scenarios. Collectively, these advancements represent a major leap towards more comprehensive and precise environmental perception in autonomous driving, addressing critical challenges and enhancing driving safety.
II-B Federated Learning
Federated learning, initially proposed as FedAvg by [37], marks a significant shift in machine learning by enabling distributed model training across multiple systems without accessing users’ raw data, thus enhancing data privacy. A primary challenge in federated learning is the management of data heterogeneity and distribution shifts, which complicate efficient model optimization. The work [38] introduces a method to mitigate client heterogeneity by incorporating proximal terms, ensuring that local model updates align with the global model. Additionally, the work [39] utilizes knowledge distillation techniques to aggregate locally computed logits, allowing for the synthesis of cohesive global models without requiring uniform local model architectures. Further addressing this heterogeneity, the works [40, 38] propose algorithms that adjust the training objective to tackle these challenges effectively. The work [41] introduced SCAFFOLD, which employs control variables for variance reduction to correct ‘client-drift’ in local updates. Moreover, the work [42] presents FedSAM, an algorithm that integrates sharpness aware minimization (SAM) as a local optimizer with a momentum-based federated learning algorithm to enhance synergy between local and global models.
II-C Federated Learning with Cooperative Perception
In CP involving CAVs and RSUs, the exchange of sensor data raises significant privacy concerns. This data, sourced from vehicles and smart city infrastructure, can reveal sensitive information about individuals, such as location and personal habits. To address this, the work [43] applied federated learning to object detection tasks in AVs, capitalizing on the method’s ability to train models without compromising data privacy. However, the broader application of federated learning in CP was initially underexplored. Recent studies such as [22] and [44] discuss the application of federated learning in CP settings. The work [45] introduces an innovative federated learning framework, H2-Fed, designed to handle Hierarchical Heterogeneity, significantly improving conventional pre-trained deep learning models for CP. Furthermore, the work [46] proposes a new client selection pipeline using V2X communications, optimizing client selection based on predicted communication latency in real CP scenarios. These developments underscore the potential for federated learning for collaborative model training in privacy-sensitive environments. Despite these advances, implementing CP through federated learning presents challenges, primarily due to heterogeneity in data distribution, computational resources, and network connectivity among devices. This variability can significantly affect learning efficiency and effectiveness. To address these challenges, the work [47] introduces a fair and efficient algorithm to manage the imbalanced distribution of data and fluctuating channel conditions. Moreover, the work [27] proposed FedBEVT, a federated transformer learning approach that focuses on BEV perception to tackle common data heterogeneity issues in CP.
In this paper, we introduce an innovative federated learning framework tailored for V2X networks in CP, specifically addressing challenges posed by heterogeneity. Our framework is distinguished by its focus on the heterogeneity caused by Non-IID data across different traffic agents. We propose a novel loss function and an advanced aggregation method designed to effectively manage the individual heterogeneity inherent in different agents, thereby enhancing the overall efficacy of model training and addressing the unique challenges of Non-IID data distribution.
III Methodology
III-A Problem Formulation
In this paper, we establish a federated learning network consisting of $M$ clients, each equipped with a unique label $m\in\{1,2,...,M\}$. The $m$-th client contains a set of data samples denoted by $\mathbf{N}_{m}=\{(\mathbf{N}_{m,i},\mathbf{Y}_{m,i})\}_{i\in\{1,2,...,|\mathbf{N}_{m}|\}}$, where ||$\cdot$|| ⋅ | denotes the number of elements in the set, $\mathbf{N}_{m,i}$ is the set of images in the $i$-th data sample captured from a multi-view camera system, and $\mathbf{Y}_{m,i}$ is the set of corresponding BEV semantic masks. Within this framework, each client independently trains a local model using their dataset and then transmits the model updates to a central server. The server integrates these updates to improve the global model. This cycle of local training and central updating repeats across several communication rounds to refine the global model optimally.
The primary optimization objective in the federated learning framework is encapsulated by:
$L=\frac{1}{\sum_{m=1}^{M}|\mathbf{N}_{m}|}\sum_{m=1}^{M}\sum_{i=1}^{|\mathbf{N}_{m}|}L_{i},$ (1)
where $L_{i}$ denotes the loss function for the data sample, and $|\mathbf{N}_{m}|$ is the number of data samples per client. This loss function is crucial for assessing and minimizing the error throughout the distributed network of clients.
Building on the framework from [27], our model features a transformer-based BEV architecture. It comprises an image feature encoder utilizing a CNN, a BEV transformer with a multi-layer attention structure (incorporating both sparse cross-view and self-attention mechanisms), and a BEV decoder. This configuration effectively encodes images from multiple cameras and transforms them into BEV features. Processed through FedBEVT, these features are decoded to produce semantic segmentation outputs from the BEV perspective, which are essential for accurately determining the positions of surrounding vehicles on the motion plane.
Our framework introduces a novel aggregation method and a customized loss function designed to tackle two primary challenges: (i) the heterogeneity of data and distribution shifts across different clients, and (ii) the varying contributions of different clients to the convergence of the global model. These enhancements are vital to ensure robust and precise model performance in autonomous driving environments.
III-B Federated Dynamic Weighted Aggregation
To address the specific challenges of applying federated learning in CP, particularly client drift due to diverse sensor types and Non-IID data distributions, we introduce the federated dynamic weighted aggregation (FedDWA) algorithm. This algorithm updates the client model described by:
$G_{m}=G_{m}^{-}-\frac{1}{|\mathbf{N}_{m}|}U_{m},$ (2)
$U_{m}=\eta_{m}(\sum_{i}^{|\mathbf{N}_{m}|}\partial_{i}(G_{m}^{-})+c_{g}^{-}-c_{m}^{-}),$ (3)
where $G_{m}$ represents the local model of the $m$-th client,$\eta_{m}$ is the local step-size, and $\partial_{i}(G_{m}^{-})$ denotes the gradient of the $i$-th data sample for the local model in the previous round. Each client uses a local control variable $c_{m}$ to account for variances in model updates, aligning local objectives with the global model to mitigate client-specific drifts, and $c_{m}^{-}$ denotes the local control variable in the previous round. Both $c_{m}$ and $c_{g}$ are initialized as 0.
Inspired by [41], we maintain a global control variable $c_{g}$ on the server, aggregating information from all clients to guide the overall direction of model updates. Diverging from [41], which uses the average of local variables to update $c_{g}$, we utilize Kullback-Leibler divergence (KLD) [48] to dynamically weigh the contributions of different clients like:
Input :
1 Number of communication round $R$ and data samples $\mathbf{N}_{m}$
2 Server input: global model $G_{g}^{-}$ and global control variable $c_{g}^{-}$ in previous round, global learning rate $\eta_{g}$
3 $m$-th client input: local control variable $c_{m}^{-}$ in previous round and local step-size $\eta_{m}$
4 for each round $r=1,\dots,R$  do
5       for  do
6
7            send $G_{g}^{-}$, $c_{g}^{-}$ to all clients $m\in\{1,2,...,M\}$
8             for  client $m\in\{1,2,...,M\}$ parallel do
9                   initialize local model on $m$-th client $G_{m}\leftarrow G_{g}^{-}$
10                   initialize $U_{m}=0$
11                   initialize $W_{m}=0$
12                   initialize $O_{m}=0$
13
14                  for each data sample $\mathbf{N_{m,i}}$ do
15
                        // Local model training
16                         calculate mini-batch gradient $\partial_{i}(G_{m})$
17                         $U_{m}$ +=+=+ = $\partial_{i}(G_{m}^{-})$
18                         $W_{m}$ +=+=+ = $\partial_{i}(G_{m})$
19                         $O_{m}$ +=+=+ = $\mathcal{D}_{\mathrm{KL}}\left(P_{g}^{-}(\mathbf{N}_{m,i}),P_{m}(\mathbf{N}_{m,i})\right)$
20                   end for
21                  $U_{m}$ $=$ $\eta_{m}(U_{m}+c_{g}^{-}-c_{m}^{-})$
                   // Local control variable update
22                   $G_{m}=G_{m}^{-}-\frac{1}{|\mathbf{N}_{m}|}U_{m}$
23                   $c_{m}=\frac{1}{|\mathbf{N}_{m}|}W_{m}$
24                   $T_{m}=\frac{1}{|\mathbf{N}_{m}|}c_{m}\cdot O_{m}$
25                   send $G_{m}$, $T_{m}$ to global agent
26
27             end for
28
29       end for
      // Global control variable update.
30
31       $c_{g}=c_{g}^{-}+\frac{1}{M}\sum_{m=1}^{M}T_{m}$
       // Global model update.
32
33       $G_{g}=G_{g}^{-}+\frac{\eta_{g}}{M}\sum_{m=1}^{M}(G_{g}^{-}-G_{m})$
34
35 end for
Algorithm 1 FedDWA
$c_{g}=c_{g}^{-}+\frac{1}{M}\sum_{m=1}^{M}T_{m},$ (4)
where $T_{m}$ is the intermediate variable, as shown in:
$T_{m}=\frac{1}{|\mathbf{N}_{m}|}c_{m}\cdot O_{m},$ (5)
and
$O_{m}=\sum_{i}^{|\mathbf{N}_{m}|}\mathcal{D}_{\mathrm{KL}}\left(P_{g}^{-}(\mathbf{N}_{m,i}),P_{m}(\mathbf{N}_{m,i})\right).$ (6)
The KLD is calculated between the predicted data distribution of the global model, $P_{g}^{-}(\mathbf{N}_{m,i})$ from the previous round, and the distribution of $m$-th local client, $P_{m}(\mathbf{N}_{m,i})$. Where $P_{g}^{-}(\mathbf{N}_{m,i})$ is approximated by using the segmentation results of our global model and $P_{m}(\mathbf{N}_{m,i})$ is approximated by using the segmentation results of our local model.
This metric helps to properly weigh the contribution of various clients to the convergence of the server model. Each client uses gradients to update the local control variable $c_{m}$:
$c_{m}=\frac{1}{|\mathbf{N}|_{m}}W_{m},$ (7)
where
$W_{m}=\sum_{i}^{|\mathbf{N}_{m}|}\partial_{i}(G_{m}).$ (8)
The global model is updated as outlined below:
$G_{g}=G_{g}^{-}+\frac{\eta_{g}}{M}\sum_{m=1}^{M}(G_{g}^{-}-G_{m}),$ (9)
where $\eta_{g}$, the learning rate, plays a crucial role in balancing the updates, and $G_{g}$ represents the global model. This equation captures the aggregation of differences between the global and local models, promoting the convergence of the global model.
III-C Dynamic Adjusting Loss Function
The cross-entropy loss function is widely used within federated learning frameworks to train individual clients. This function effectively gauges the performance of classification models that output probabilities between 0 and 1. Nevertheless, the inherent data heterogeneity across clients, particularly with non-IID data, complicates the convergence to a global optimum.
To address this issue, we introduce an enhanced local training loss function termed dynamic adjusting loss (DALoss):
$L_{i}=\mathcal{S}\left(F(\mathbf{N}_{m,i}),\mathbf{Y}_{m,i}\right)+Q_{i},$ (10)
where $\mathcal{S}\left(F(\mathbf{N}_{m,i}),\mathbf{Y}_{m,i}\right)$ denotes the standard cross-entropy loss for the $i$-th data sample of client model $F(\mathbf{N}_{m,i})$ against the true labels $\mathbf{Y}_{m,i}$, and $Q_{i}$ is the dynamic control term defined by:
$Q_{i}=C\cdot\mathcal{D}_{\mathrm{KL}}\left(P_{g}^{-}(\mathbf{N}_{m,i}),P_{m}(\mathbf{N}_{m,i})\right)\cdot||G_{g}^{-}-G_{m}||_{2}^{2},$ (11)
where $C\cdot\mathcal{D}_{\mathrm{KL}}\left(P_{g}^{-}(\mathbf{N}_{m,i}),P_{m}(\mathbf{N}_{m,i})\right)$ scales the quadratic penalty based on the KL divergence, measuring the discrepancy between the probabilistic distributions of the global and local models from the previous communication round. $C$ is a hyper-parameter, and $||\cdot||_{2}$ denotes the $L^{2}$ norm. This divergence quantifies the differences, ensuring that local updates contribute significantly toward global model convergence.
This modification incorporates a dynamic adjustable regularization mechanism to better align each client’s model updates with the global model. This dynamic adjustment of the regularization parameter allows our approach to adapt to the evolving nature of the training process and the unique characteristics of each client. By accommodating varying client impacts and data heterogeneity, our method aims to enhance both accuracy and convergence efficiency, thereby achieving a more globally optimized model in CP applications.
IV Experiments and Analysis
The experiments were conducted on a computer equipped with a single NVIDIA GeForce RTX 4090 GPU and a 12th Generation Intel Core i9-12900K CPU. The software environment consisted of Pytorch 2.0.1 and CUDA 12.3.
IV-A Experiment Setup
Our experimental evaluation leveraged the extensive and publicly available CP datasets from OPENV2V [29], which were augmented with an additional dataset as detailed in [27]. This enhancement broadens the diversity and volume of the scenarios represented, providing a robust foundation for testing our model.
TABLE I: Comparison between Our Method and Other State-of-the-art Methods.
Method Client 1: Bus Client 2: Truck Client 3: Car A Client 4: Car B
N1 = 1388 N2 = 1488 N3 = 2140 N4 = 1384
[…]
  • $\uparrow$
    The upper arrow denotes that better performance is registered with higher value.
  • $\downarrow$
    The lower arrow denotes that better performance is registered with lower value.
Following the methodology established in [27], our study orchestrates a collaborative framework among four distinct industrial clients to assess the efficacy of our proposed object detection model under various operational conditions. Specifically, the dataset comprises:
  • Client 1 Bus: 1,398 frames in the training set and 413 frames in the testing set.
  • Client 2 Truck: 1,459 frames in the training set and 356 frames in the testing set.
  • Client 3 Car A: 11,636 frames in the training set and 3,546 frames in the testing set.
  • Client 4 Car B: 7,142 frames in the training set and 10,191 frames in the testing set.
These entities collaboratively trained their models using a federated learning approach, which improved their perception capabilities. This arrangement ensured that while each entity benefited from shared data to enhance their models, direct access to the raw sensor data of others was strictly prohibited, addressing privacy concerns.
For a fair comparison between our federated learning strategy and state-of-the-art approaches, our model architecture adheres to the configuration described in [27]. This involves processing camera images observed by clients. Similar to [27], we begin by sending these images through a 3-layer ResNet34 encoder, which is proficient at extracting and encoding image features across various spatial dimensions. This encoding process yields feature tensors at differentiated spatial resolutions: $64\times 64\times 128$, $32\times 32\times 256$, and $16\times 16\times 512$.
(a) Bus client
(b) Truck client
(c) Car client A
(d) Car client B
Figure 1: IoU Results
Subsequently, we employ a focused attention cross (FAX) transformer operation, which is attention-based and skillfully manages interactions between BEV embeddings and the encoded image features, conceptualized as query, key, and value within this framework. Finally, a 3-layer bilinear module is used to transform these enriched BEV features. In our model configuration, the hyper-parameter $C$ defined in Equation (11) is set to 0.1. For other hyper-parameters, we adopt the configurations detailed in [27]. We selected these settings due to their proven effectiveness in similar contexts, ensuring that our model benefits from established best practices while maintaining consistency with state-of-the-art for fair comparison.
IV-B Performance Analysis
In our comparative study, we evaluate the performance of our proposed algorithm, FedDWA, against several established methods, including FedAvg [37], FedRep [49], FedTP [50], and FedBEVT [27]. The evaluation focuses on the average intersection over union (IoU) achieved by models across three use cases (UCs) and the number of communication rounds required to attain optimal IoU values.
The experimental analysis presented in Table I provides a comprehensive comparison of our federated learning framework against these prominent approaches.
Performance metrics include training loss, IoU, and the number of communication rounds (Com.) required to achieve peak IoU, offering a detailed assessment of the efficacy and efficiency of our approach within a federated learning context.
The key findings for each client scenario are as follows:
[…]
Fig. 1 illustrates the relationship between the number of communication rounds and IoU during the testing phase for the four clients, providing a detailed view of test progression. Notably, the truck client and car client B reached optimal IoU with fewer communication rounds, potentially due to data distributions more closely aligned with the predictions of their local models. However, after reaching peak IoU, there is a noticeable decline in performance, suggesting possible overfitting as the models become excessively tailored to their local datasets, to the detriment of generalization.
IV-C Ablation Experiment
TABLE II: Ablation Experiment for Effectiveness of DALoss with IoU
Method Client 1 Client 2 Client 3 Client 4
[…]
V Conclusion
In this paper, we explore the effectiveness of federated learning frameworks for optimizing transformer-based models in BEV perception within road traffic datasets. A primary focus of our investigation is the challenge of data heterogeneity, which significantly impedes the performance of federated learning methodologies. To address this issue, we introduce an innovative algorithm, FedDWA, together with DALoss, specifically designed to mitigate the adverse effects of data diversity. Our comparative analyses with existing federated learning frameworks demonstrate that our proposed solutions significantly enhance the efficacy of the federated learning framework. These improvements are primarily achieved through dynamic adjustments of the model convergence direction and the equitable integration of diverse local models into a unified global model. Our findings underscore the viability and potential of federated learning frameworks in BEV perception, suggesting the approach as a resilient, efficient, and inclusive alternative for model training.